OpenStudy (amy0799):

verify the identity cot(x-pi/2)=-tan x

OpenStudy (amy0799):

one side

OpenStudy (amy0799):

oh it's fine, take your time, im just glad ur helping

OpenStudy (solomonzelman):

I think I am going the wrong direction. I haven't done these for a hundred years...

OpenStudy (freckles):

so cot is and odd function that is cot(-x)=-cot(x) so we have that cot(x-pi/2)=-cot(pi/2-x) i love to draw triangles...

OpenStudy (freckles):

 the other angle is pi/2-x

OpenStudy (freckles):

find tan(x) and cot(pi/2-x)

OpenStudy (freckles):

and compare the values

OpenStudy (amy0799):

oh ok, thank you for trying

OpenStudy (freckles):

that is my favorite way to pick when doing cofunctions

OpenStudy (solomonzelman):

it's cot(x-pi/2)=-tan x :)

OpenStudy (freckles):

but you could also expand the cot(x-pi/2) using other identities

OpenStudy (freckles):

that is correct @SolomonZelman but cot(x-pi/2)=-cot(pi/2-x)

OpenStudy (freckles):

since cot is odd

OpenStudy (solomonzelman):

yes

OpenStudy (freckles):

oh i thought you were saying something about what i was saying

OpenStudy (amy0799):

how do you get one side to equal the other side?

OpenStudy (freckles):

You can prove the identity by using the fact that cot is odd and then construct a right triangle if you didn't know that the following is a co-function identity cot(pi/2-x)=tan(x)

OpenStudy (amy0799):

so the answer would be because it's a cofunction identity and it's odd?

OpenStudy (freckles):

cot(x-pi/2)=cot(-(pi/2-x))=-cot(pi/2-x) since cot(x) is an odd function cot(pi/2-x)=tan(x) by co function identity -cot(pi/2-x)=-tan(x) honestly I love to show the cofunction identity but if you have already shown them in class then we don't really need to do it over again

OpenStudy (amy0799):

on the last step, the two equations aren't equal to each other

OpenStudy (freckles):

how so

OpenStudy (freckles):

if 5=5 then -5=-5

OpenStudy (amy0799):

after -cot(pi/2-x) = -tan x, would i write -tan x = -tan x?

OpenStudy (freckles):

cot(pi/2-x)=tan(x) so yeah

OpenStudy (amy0799):

ok thank you so much :D

OpenStudy (freckles):

you could also go the long way about this \[\cot(x-\frac{\pi}{2})=\frac{\cos(x-\frac{\pi}{2})}{\sin(x-\frac{\pi}{2})} \\ =\frac{\cos(x)\cos(\frac{\pi}{2})+\sin(x)\sin(\frac{\pi}{2})}{\sin(x)\cos(\frac{\pi}{2})-\sin(\frac{\pi}{2})\cos(x)} \\ =\frac{\cos(x) \cdot 0+\sin(x) \cdot 1}{\sin(x) \cdot 0 -1 \cdot \cos(x) }\]

OpenStudy (freckles):

i'm sure you can finish this way